OpenStudy (anonymous):

Use work and energy to find the average friction force that stops a bullet with a mass of 7.8kg, v1=575 m/s, x=5.5cm.

OpenStudy (vijay):

81545.4545 N (Newton )

OpenStudy (anonymous):

The answer in the back of the book is 2.34x10^4N. I got Wnet=KEf-KEi=Average Friction. I got \[-FkDeltax = 1/2mVf ^{2}-1/2mVi ^{2}\]. Then, I isolated -Fk. But I don't know how to find the Vf.